04 Question (2 points) See page 78 In addition to finding the optimal bundles given prices and income, utility maximization cd
Answer #1

u(x,y) = 100xy

\partialu/\partialx = MUx = 100y

\partialu/\partialy = MUy = 100x

MRSx,y   = MUx/MUy

= 100y/100x

= y/x

Budget constraint

Pxx + Pyy = m (i)

At optimal choice

MRSx,y = Px/Py

y/x = Px/Py

y = (Px/Py)x

substitute    y = (Px/Py)x in equation (i)

Pxx + Pyy = m

Pxx + Py(Px/Py)x = m  

Pxx + Pxx = m

2Pxx = m

x = m/2Px  

x*(Px,Py,m) =   m/2Px  

y = (Px/Py)x

=    (Px/Py)( m/2Px )

= m/2Py  

y*(Px,Py,m) =   m/2Py